Gas in a piston-cylinder apparatus is heated with 150.0 J while simultaneously decreasing its volume from 1.500 L to 0.500 L at a constant pressure of 1.000 bar. Calculate the total energy change in Joules for this process.

Options:

A

0 (the work and heat contributions cancel out)

B

-150.0 J

C

-50.0 J

D

+50.0 J

E

+250.0 J
Answer #1

W = -Pext(Vf-Vi)
P= 1.0 bar
= (1.0/1.01325) atm
= 0.9869 atm
Vi = 1.5 L
Vf = 0.5 L
put values in above expression
W = -Pext(Vf-Vi)
W = -0.9869 atm (0.5L - 1.5L)
W = 0.9869 atm.L
W = 0.9869 *101.33 J
W = 100.00 J

Given:
Q = 150.0 J

use:
ΔU = Q + W
ΔU = (150.0 J) + (100.00 J)
ΔU = 250.0 J
Answer: E